OpenStudy (anonymous):

2-Uranium has several isotopes , one of the them has mass number of 235 and atomic mass of 92 ,calculate number of neutrons and state the number of electrons.

OpenStudy (anonymous):

no of electrons = atomic no in a neutral atom that is in this question. so electons = 92. also protons = neutrons in neutral atom so protons are also 92. we know that mass no = neutrons +protons so 235= n+92. hence n=235-92 n=143

OpenStudy (anonymous):

im sorry Protons = no of electrons in neutral atom ** .... so protons are also 92